Hi! Can anyone help me discern the solution and answer for this problem:
"The supply equation for a calculator is y=m^2+ square root of 'm', where 100y calculators are supplied when m dollars is the price per calculator. find the instantaneous rate of change of the supply with respect to the price when the price is $1,600."
Thanks! :)

Answer by stanbon(75887)

"The supply equation for a calculator is y=m^2+ square root of 'm', where 100y calculators are supplied when m dollars is the price per calculator. find the instantaneous rate of change of the supply with respect to the price when the price is $1,600."
-----
The instantaneous rate of change = the derivative is y'(m) = 2m+(1/(sqrt(m))
---
y'(1600) = 2*1600+(1/40) = 3200.03
==================
Cheers,
Stan H.
==================